1. Calculate the molecular masses of

A. C2H2

B. N2H4

C. Na2O

D. O3

E. pH3

F. CO2

G. C3HgO


2. Calculate the density of the following

A. 93.6g occupying 15mL

B. 0.992g occupying 1,005mL

C. 13.7g occupying 11.4mL

D. 16.8g occupying 3.19mL.



1
Expert's answer
2021-10-18T06:49:11-0400

1) A. (12×2)+ 2= 26

B. 14× 2+ 4 = 32

C. 23× 2 + 16= 62

D. 16 × 3 = 48

E. 31+ 3= 34

F. 12+ 32= 44

G.12 × 3 + 9 + 16= 61

2) density, "D = M\/ v"

A. 93.6/ 15= 5.85g/cm3

B. 0.993/ 1005= 0.000985 g/cm3

C. 13.7/11.4= 1.2017g/cm3

D.16.8/ 3.19 = 5.264g/cm3